# A Survey of WebAgents: Towards Next-Generation AI Agents for Web Automation with Large Foundation Models

### v2

@@ 1 Introduction
To fully explore the potential of LFMs, recent efforts have been made to advance LFMs-empowered WebAgents to complete various web tasks according to user instructions (Hu et al., 2024). For instance, the recent debut of a novel AI Agent framework named AutoGPT has attracted significant interest from both academic and industrial communities, which exhibits impressive capabilities in autonomously handling complex tasks across both work and daily environments (Significant Gravitas, [n. d.]). Unlike chatbots, AutoGPT can plan and execute multimodal tasks independently, performing automated searches and multi-step actions without requiring ongoing user instructions and supervision. In this context, as illustrated in Figure 1, users only need to provide a natural language instruction, such as ‘Schedule a meeting with Leon at Starbucks on November 23, 2024, at 4:00 pm via email.’ WebAgents can autonomously open the ‘Email’ application, retrieve Leon’s email address, compose the email, and send it, thereby automating the entire scheduling process and greatly enhancing the convenience of daily life. Given the remarkable progress in developing LFM-empowered WebAgents and the growing number of related studies, there is a pressing need for a systematic review of recent advances in this field.
To bridge this gap, this survey provides a comprehensive overview of WebAgents by summarizing representative methods from the perspectives of architecture, training, and trustworthiness. Specifically, Section 2 introduces the background knowledge of traditional agents and the fundamental pipeline of WebAgents. In Section 3, we review existing studies based on the three processes of WebAgents: perception, planning and reasoning, and execution. Next, we summarize two crucial aspects (i.e., data and training strategies) in the training of WebAgents in Section 4. After that, we review studies that focus on investigating the trustworthy WebAgents, including their safety, robustness, privacy, and generalizability, in Section 5. Finally, in Section 6, we discuss promising future research directions in WebAgents.
@@ 3 WebAgent Architectures
There are three crucial and consecutive processes for WebAgents to fulfill user commands: 1) Perception requires WebAgents to accurately observe the current environment, 2) Planning and Reasoning require WebAgents to analyze based on the current environment, interpret user-given tasks, and predict reasonable next actions, and 3) Execution requires that WebAgents perform the generated actions and interact with the environment effectively. In the following section, we will comprehensively review the important techniques employed by WebAgents during these processes, and the details are also summarised in Table 1.
@@ 3.1. Perception
With the advancement of large language models, extensive studies have been proposed to leverage its human-like understanding and reasoning abilities to assist users in addressing complex tasks. Since LLMs can only handle natural language, these WebAgents usually leverage the textual metadata of webpages (e.g., HTML and accessibility trees) to perceive the environment (Kim et al., 2023; Lo et al., 2023; Zheng et al., [n. d.]; Ma et al., [n. d.]). For example, MindAct (Deng et al., 2023) introduces a two-stage framework that combines a fine-tuned small LM with an LLM to efficiently process large HTML documents, significantly reducing the input size while preserving essential information. This approach enables accurate prediction of both the target element and the corresponding action, effectively balancing efficiency and performance in web-based tasks. Gur et al. (2024b) introduce an LLM-driven agent that learns from self-experience to complete tasks on real-world webpages. It summarizes long HTML documents into task-relevant snippets to extract the environmental information and decomposes user instructions into sub-tasks for effective planning.
@@
Despite the remarkable success of text-based WebAgents, leveraging the textual metadata of the environment usually fails to align closely with human cognitive processes since the Graphical User Interfaces (GUI) are inherently visual (Shen et al., 2024b; Xu et al., 2024d). Additionally, textual representations usually vary across different environments and are verbose, leading to poor generalization abilities and increased computational overhead (Zhang and Zhang, 2024b). Recently, breakthroughs in large vision-language models (VLMs) have significantly enhanced the capabilities of AI systems in processing complex visual interfaces. To leverage the visual understanding capabilities of VLMs, numerous studies have integrated them into WebAgents, utilizing screenshots to perceive the environment (Zhang and Zhang, 2024b; Guan et al., 2023; Gao et al., 2023; Kil et al., 2024). SeeClick (Cheng et al., 2024) only relies on screenshots as observations to predict the next action and enhances the agent’s ability to locate relevant visual elements within screenshots by introducing a grounding pre-training process. OmniParser (Lu et al., 2024) introduces an effective method to parse user interface screenshots into structured elements and enhances GPT-4V’s ability to accurately ground actions to specific regions on the screen. OmniParser employs specialized models for parsing interactable regions on the screen and capturing the functions of detected elements, significantly improving the agent’s performance on the visual benchmark.
@@
In addition to solely utilizing textual metadata or screenshots to comprehend the environment, numerous studies also leverage multi-modal data, combining their complementary strengths to provide WebAgents with a more comprehensive environmental perception (Wang and Liu, 2024; Iong et al., 2024; Lai et al., 2024). For instance, MMAC-Copilot (Song et al., 2024a) integrates GPT-4V (Yang et al., 2023a) for interpreting visual information from screenshots while leveraging Gemini Vision (Li et al., 2024d) to process and analyze video content, significantly enhancing the model’s capabilities in handling multi-modal data. WebVoyager (He et al., 2024) is a multi-modal WebAgent that autonomously completes web tasks end-to-end by processing both screenshots and textual content from interactive web elements. It leverages Set-of-Mark prompting (Yang et al., 2023b) to overlay bounding boxes of the interactive elements on the webpages, significantly enhancing the agent’s decision-making ability and enabling accurate action prediction and execution.
3.2. Planning and Reasoning
@@
Implicit planning WebAgents directly feed the user instruction and the environmental observation into agents without an explicit task decomposition process (Ma et al., [n. d.]; Cho et al., 2024; Koh et al., 2024). For instance, WebWISE (Tao et al., 2023) directly provides the task message to the agent and prompts it to generate actions progressively based on filtered Document Object Model (DOM) elements as environmental observations. OpenWebAgent (Iong et al., 2024) first uses a Web Processing Module to extract and simplify HTML elements, perform Optical Character Recognition (OCR) (Singh et al., 2012) on screenshots, and annotate interactive elements. Then, OpenWebAgent introduces an Action Generation Module to predict the next action by directly feeding the current task and webpage context into agents.
@@ 4.1. Data
Data collection involves gathering data from public datasets or directly from real-world scenarios. For instance, Lexi (Banerjee et al., 2023) collects and curates 114k UI images paired with functional captions sourced from open websites, exhibiting remarkable diversity across applications, platforms, and UI characteristics. Rather than indiscriminately aggregating all available data, some studies like ShowUI (Lin et al., 2024) demonstrate that cautiously sampling high-quality, representative data from public datasets can significantly enhance WebAgents’ performance. While data annotated by human experts exhibits high quality, its labor-intensive process limits the total amount of the resulting datasets, leading to undertrained models with limited generalization to unseen online contexts. To address this challenge, several automated generation methods have been proposed to construct cost-effective datasets without reliance on human assistance or large vision-language models (Banerjee et al., 2023; Lee et al., 2025). For instance, Falcon-UI (Shen et al., 2024b) autonomously generates a dataset, Insight-UI, by interacting with publicly available webpages, comprising multi-step, cross-platform screenshots with all visible interactive elements and character-level OCR annotations. ScribeAgent (Shen et al., 2024a) collects action sequence data across various websites executed by real users from Scribe. These sequences are structured as document-based workflows, encompassing a diverse range of targets and difficulty levels. Besides, UINav (Li et al., 2023a) mitigates the pressure of scarce training data by randomizing the attributes of secondary UI elements, such as text embedding or element offsets, to augment demonstrations without additional data collection overhead.
@@ 4.2. Training Strategies
The rapid evolution of LFMs, endowed with human-like intelligence and robust comprehension of visual and textual information, substantially boosts the development of intelligent WebAgents. Building upon these capabilities, training-free methods directly adapt LFMs as specialized WebAgents, harnessing well-crafted prompts to guide models in executing web tasks without any architectural modifications or parameter updates (Chen et al., 2024a; Rawles et al., 2023; Wei et al., 2022; Wen et al., 2023). For example, Auto-GUI (Zhang and Zhang, 2024a) synthesizes both previous action histories and future action plans within a chain-format prompt to inform the decision-making process for determining the subsequent action step. Advancing this methodology, CoAT (Zhang et al., 2024d) proposes a chain-of-action-thought prompting paradigm, which integrates actions and thoughts to enable more efficient navigation capabilities. The paradigm incorporates screen descriptions, previous actions, and their outcomes to infer the reasoning behind the next action determination, alongside generating textual descriptions of the subsequent step and its potential results.
@@
Although general-purpose LFMs are extensively trained on enormous datasets, these models often fall short in GUI-aware capabilities, particularly screen understanding and Optical Character Recognition (OCR), hindering them to comprehend and interact effectively with webpages. Specifically, general LFMs might focus on decorative icons or background text instead of the key interface elements, leading to misinterpretations of available functionalities (Liu et al., 2024c). To bridge this gap, numerous GUI comprehension training approaches are developed to enhance the critical foundational GUI understanding capabilities of WebAgents through supervised learning on large-scale web datasets (Shen et al., 2024b; Wu et al., 2024d; Dorka et al., 2024; Banerjee et al., 2023). For instance, Aguvis (Xu et al., 2024d) introduces a two-stage training paradigm that begins with unifying GUI environments as images and concentrates on equipping the model to comprehend and interact with objects within a single GUI screenshot during the pre-training stage, laying a robust GUI comprehension foundation for subsequent fine-tuning. Similarly, OS-Atlas (Wu et al., 2024c) leverages vast and diverse high-quality triplets ¡screenshot, referring expression, coordinates¿ to train an existing VLM by predicting the element coordinates with screenshots and referring expression, which enables the model to effectively interpret GUI screenshots and accurately identify their constituent elements and locations. While Aguvis and OS-Atlas focus primarily on vision-centric GUI understanding learning, MM1.5 (Zhang et al., 2024a) capitalizes on text-rich OCR data during training to significantly enhance its ability to comprehend interleaved image-text data. This capability is highly required in text-rich scenarios like E-commerce platforms, where product webpages often contain a large amount of detailed text descriptions, such as product specifications and reviews, as well as numerous images, like product photos and size charts. In addition, given the context sensitivity of the UI elements, LVG (Qian et al., 2024) introduces layout-guided contrastive learning, which captures the semantics of individual UI elements based on their visual organization, further enhancing the model’s GUI comprehension.
@@
Although WebAgents can effectively understand environmental information through GUI comprehension training, accurately reasoning and generating the next steps for web interactions based on user tasks remains a significant challenge due to the complexity of web environments and the diversity of user objectives in web automation. Therefore, task-specific fine-tuning aims to equip WebAgents with web task-oriented skills, such as planning, reasoning, and interacting capabilities (Shen et al., 2024b; Furuta et al., 2023b; Xu et al., 2024d; Guan et al., 2023). For instance, Gur et al. (2023) propose an LLM-driven agent fine-tuned with scripted planning datasets, which decomposes natural language instructions into manageable sub-instructions, summarizes lengthy HTML documents into task-relevant snippets, and executes actions through self-generated Python programming. LCoW (Lee et al., 2025) enhances WebAgents’ decision-making capabilities by leveraging contextualized observations from a Contextualization Module. Specifically, this module is iteratively fine-tuned by the optimal contextualized observations sampling from its own outputs. Besides, to efficiently navigate in the exponentially large exploration space, NNetNav (Murty et al., 2024) utilizes the hierarchical structure of language instructions for fine-tuning, rendering the search process more manageable. It prunes interaction episodes automatically when an intermediate trajectory cannot be meaningfully annotated, optimizing exploration effectiveness. WebGUM (Furuta et al., 2023a) transforms web navigation tasks into the instruction-following task that enables WebAgents to perform a data-driven offline supervised fine-tuning paradigm with instruction-following format problems and chain-of-thought examples across a variety of domains. The web navigation action output is in free-form text, enhancing the flexibility and adaptability to tackle real-world web navigation. In addition, LLMPA (Guan et al., 2023) implements an end-to-end finetuning framework to accomplish multi-step operations based on high-level user command by integrating various auxiliary modules, like instruction decomposition and action prediction. Notably, apart from these key supporting modules, LLMPA employs Controllable Calibration to meticulously evaluate predicted actions, effectively mitigating the hallucination issues prevalent in LLMs and enhancing overall execution reliability.
@@ 5 Trustworthy WebAgents
In recent years, the rapid evolution of LFM-empowered agents has ushered in a new era, profoundly reshaping various facets of web applications, including E-commerce, healthcare, and education. However, alongside this remarkable progress, a growing body of research and real-world incidents has highlighted the inherent risks and challenges associated with WebAgents (Yang et al., 2025; Schwartz et al., 2023; Deng et al., 2024). Recent studies have begun to reveal the potential threats posed by agents closely integrated with web systems. These threats range from unreliable and opaque decision-making in safety-critical contexts to the perpetuation of bias and unfairness towards marginalized communities (Deng et al., 2024; Li et al., 2024a). Moreover, concerns have been raised about the inadvertent exposure of user privacy and sensitive business information, further emphasizing the critical need to ensure the security of WebAgent technologies. Additionally, limitations in generalizability pose serious risks for WebAgents when confronted with diverse and unforeseen situations, potentially leading to critical failures in handling out-of-distribution data or operating across different domains. Consequently, developing advanced techniques for trustworthy WebAgents has become an area of increasing interest for researchers, primarily focusing on creating and deploying artificial intelligence systems that prioritize Safety&Robustness, Privacy, and Generalizability.

### v3

@@ 1 Introduction
To fully explore the potential of LFMs, recent efforts have been made to advance LFMs-empowered WebAgents to complete various web tasks according to user instructions (Hu et al., 2024). For instance, the recent debut of a novel AI Agent framework named AutoGPT has attracted significant interest from both academic and industrial communities, which exhibits impressive capabilities in autonomously handling complex tasks across both work and daily environments (Significant Gravitas, [n. d.]). Unlike chatbots, AutoGPT can plan and execute complex tasks independently, performing automated searches and multi-step actions without requiring ongoing user instructions and supervision. In this context, as illustrated in Figure 1, users only need to provide a natural language instruction, such as ‘Schedule a meeting with Leon at Starbucks on November 23, 2024, at 4:00 pm via email.’ WebAgents can autonomously open the ‘Email’ application, retrieve Leon’s email address, compose the email, and send it, thereby automating the entire scheduling process and greatly enhancing the convenience of daily life. Given the remarkable progress in developing LFM-empowered WebAgents and the growing number of related studies, there is a pressing need for a systematic review of recent advances in this field.
To bridge this gap, this survey provides a comprehensive overview of WebAgents by summarizing representative methods from the perspectives of architecture, training, and trustworthiness. Specifically, Section 2 introduces the background knowledge of traditional agents and the fundamental pipeline of WebAgents. In Section 3, we review existing studies based on the three processes of WebAgents: perception, planning & reasoning, and execution. Next, we summarize two crucial aspects (i.e., data and training strategies) in the training of WebAgents in Section 4. After that, we review studies that focus on investigating the trustworthy WebAgents, including their safety & robustness, privacy, and generalizability, in Section 5. Finally, in Section 6, we discuss promising future research directions in WebAgents.
@@ 3 WebAgent Architectures
There are three crucial and consecutive processes for WebAgents to fulfill user commands: 1) Perception requires WebAgents to accurately observe the current environment, 2) Planning & Reasoning require WebAgents to analyze based on the current environment, interpret user-given tasks, and predict reasonable next actions, and 3) Execution requires that WebAgents perform the generated actions and interact with the environment effectively. In the following section, we will comprehensively review the important techniques employed by WebAgents during these processes, and the details are also summarised in Table 1.
@@ 3.1. Perception
With the advancement of large language models, extensive studies have been proposed to leverage its human-like understanding and reasoning abilities to assist users in addressing complex tasks. Since LLMs can only handle natural language, these WebAgents usually leverage the textual metadata of webpages (e.g., HTML and accessibility trees) to perceive the environment (Kim et al., 2023; Lo et al., 2023; Zheng et al., [n. d.]; Ma et al., [n. d.]). For example, MindAct (Deng et al., 2023) introduces a two-stage framework that combines a fine-tuned small language model (LM) with an LLM to efficiently process large HTML documents, significantly reducing the input size while preserving essential information. This approach enables accurate prediction of both the target element and the corresponding action, effectively balancing efficiency and performance in web-based tasks. Gur et al. (2024b) introduce an LLM-driven agent that learns from self-experience to complete tasks on real-world webpages. It summarizes long HTML documents into task-relevant snippets to extract the environmental information and decomposes user instructions into sub-tasks for effective planning.
@@
Despite the remarkable success of text-based WebAgents, leveraging the textual metadata of the environment usually fails to align closely with human cognitive processes since the Graphical User Interfaces (GUI) are inherently visual (Shen et al., 2024b; Xu et al., 2024d). Additionally, textual representations usually vary across different environments and are verbose, leading to poor generalization abilities and increased computational overhead (Zhang and Zhang, 2024b). Recently, breakthroughs in large vision-language models (VLMs) have significantly enhanced the capabilities of AI systems in processing complex visual interfaces. To leverage the visual understanding capabilities of VLMs, numerous studies have integrated them into WebAgents, utilizing screenshots to perceive the environment (Zhang and Zhang, 2024b; Guan et al., 2023; Gao et al., 2023; Kil et al., 2024). SeeClick (Cheng et al., 2024) only relies on screenshots as observations to predict the next action and enhances the agent’s ability to locate relevant visual elements within screenshots by introducing a grounding pre-training process. OmniParser (Lu et al., 2024) introduces an effective method to parse user interface screenshots into structured elements and enhances GPT-4V’s (Yang et al., 2023a) ability to accurately ground actions to specific regions on the screen. OmniParser employs specialized models for parsing interactable regions on the screen and capturing the functions of detected elements, significantly improving the agent’s performance on the visual benchmark.
@@
In addition to solely utilizing textual metadata or screenshots to comprehend the environment, numerous studies also leverage multi-modal data, combining their complementary strengths to provide WebAgents with a more comprehensive environmental perception (Wang and Liu, 2024; Iong et al., 2024; Lai et al., 2024). For instance, MMAC-Copilot (Song et al., 2024a) integrates GPT-4V for interpreting visual information from screenshots while leveraging Gemini Vision (Li et al., 2024d) to process and analyze video content, significantly enhancing the model’s capabilities in handling multi-modal data. WebVoyager (He et al., 2024) is a multi-modal WebAgent that autonomously completes web tasks end-to-end by processing both screenshots and textual content from interactive web elements. It leverages Set-of-Mark prompting (Yang et al., 2023b) to overlay bounding boxes of the interactive elements on the webpages, significantly enhancing the agent’s decision-making ability and enabling accurate action prediction and execution.
3.2. Planning & Reasoning
@@
Implicit planning WebAgents directly feed the user instruction and the environmental observation into agents without an explicit task decomposition process (Ma et al., [n. d.]; Cho et al., 2024; Koh et al., 2024b). For instance, WebWISE (Tao et al., 2023) directly provides the task message to the agent and prompts it to generate actions progressively based on filtered Document Object Model (DOM) elements as environmental observations. OpenWebAgent (Iong et al., 2024) first uses a Web Processing Module to extract and simplify HTML elements, perform Optical Character Recognition (OCR) (Singh et al., 2012) on screenshots, and annotate interactive elements. Then, OpenWebAgent introduces an Action Generation Module to predict the next action by directly feeding the current task and webpage context into agents.
@@ 4.1. Data
Data collection involves gathering data from public datasets or directly from real-world scenarios. For instance, Lexi (Banerjee et al., 2023) collects and curates 114k UI images paired with functional captions sourced from open websites, exhibiting remarkable diversity across applications, platforms, and UI characteristics. Rather than indiscriminately aggregating all available data, some studies like ShowUI (Lin et al., 2024) demonstrate that cautiously sampling high-quality, representative data from public datasets can significantly enhance WebAgents’ performance. While data annotated by human experts exhibits high quality, its labor-intensive process limits the total amount of the resulting datasets, leading to undertrained models with limited generalization to unseen online contexts. To address this challenge, several autonomous generation methods have been proposed to construct cost-effective datasets without reliance on human assistance or large vision-language models (Banerjee et al., 2023; Lee et al., 2025). For instance, Falcon-UI (Shen et al., 2024b) autonomously generates a dataset, Insight-UI, by interacting with publicly available webpages, comprising multi-step, cross-platform screenshots with all visible interactive elements and character-level OCR annotations. ScribeAgent (Shen et al., 2024a) collects action sequence data across various websites executed by real users from Scribe. These sequences are structured as document-based workflows, encompassing a diverse range of targets and difficulty levels. Besides, UINav (Li et al., 2023a) mitigates the pressure of scarce training data by randomizing the attributes of secondary UI elements, such as text embedding or element offsets, to augment demonstrations without additional data collection overhead.
@@ 4.2. Training Strategies
The rapid evolution of LFMs, endowed with human-like intelligence and robust comprehension of visual and textual information, substantially boosts the development of intelligent WebAgents. Building upon these capabilities, training-free methods directly adapt LFMs as specialized WebAgents, harnessing well-crafted prompts to guide models in executing web tasks without any architectural modifications or parameter updates (Chen et al., 2024a; Rawles et al., 2023; Wei et al., 2022; Wen et al., 2023). For example, Auto-GUI (Zhang and Zhang, 2024a) synthesizes both previous action histories and future action plans within a chain-format prompt to inform the decision-making process for determining the subsequent action step. Advancing this methodology, CoAT (Zhang et al., 2024d) proposes a chain-of-action-thought prompting paradigm, which integrates actions and thoughts to enable more efficient navigation capabilities. The paradigm incorporates screen descriptions, previous actions, and their current outcomes to offer explicit explanations behind the next action determination, alongside generating textual descriptions of the subsequent step and its potential results.
@@
Although general-purpose LFMs are extensively trained on enormous datasets, these models often fall short in GUI-aware capabilities, particularly screen understanding and OCR, hindering them to comprehend and interact effectively with webpages. Specifically, general LFMs might focus on decorative icons or background text instead of the key interface elements, leading to misinterpretations of available functionalities (Liu et al., 2024c). To bridge this gap, numerous GUI comprehension training approaches are developed to enhance the critical foundational GUI understanding capabilities of WebAgents through supervised learning on large-scale web datasets (Shen et al., 2024b; Wu et al., 2024d; Dorka et al., 2024; Banerjee et al., 2023). For instance, Aguvis (Xu et al., 2024d) introduces a two-stage training paradigm that begins with unifying GUI environments as images and concentrates on equipping the model to comprehend and interact with objects within a single GUI screenshot during the pre-training stage, laying a robust GUI comprehension foundation for subsequent fine-tuning. Similarly, OS-Atlas (Wu et al., 2024c) leverages vast and diverse high-quality triplets ¡screenshot, referring expression, coordinates¿ to train an existing VLM by predicting the element coordinates with screenshots and referring expression, which enables the model to effectively interpret GUI screenshots and accurately identify their constituent elements and locations. While Aguvis and OS-Atlas focus primarily on vision-centric GUI understanding learning, MM1.5 (Zhang et al., 2024a) capitalizes on text-rich OCR data during training to significantly enhance its ability to comprehend interleaved image-text data. This capability is highly required in text-rich scenarios like E-commerce platforms, where product webpages often contain a large amount of detailed text descriptions, such as product specifications and reviews, as well as numerous images, like product photos and size charts. In addition, given the context sensitivity of the UI elements, LVG (Qian et al., 2024) introduces layout-guided contrastive learning, which captures the semantics of individual UI elements based on their visual organization, further enhancing the model’s GUI comprehension.
@@
Although WebAgents can effectively understand environmental information through GUI comprehension training, accurately reasoning and generating the next steps for web interactions based on user tasks remains a significant challenge due to the complexity of web environments and the diversity of user objectives in web automation. Therefore, task-specific fine-tuning aims to equip WebAgents with web task-oriented skills, such as planning, reasoning, and interacting capabilities (Shen et al., 2024b; Furuta et al., 2023b; Xu et al., 2024d; Guan et al., 2023). For instance, Gur et al. (2023) propose an LLM-driven agent fine-tuned with scripted planning datasets, which decomposes natural language instructions into manageable sub-instructions, summarizes lengthy HTML documents into task-relevant snippets, and executes actions through self-generated Python programming. LCoW (Lee et al., 2025) enhances WebAgents’ decision-making capabilities by leveraging contextualized observations from a Contextualization Module. Specifically, this module is iteratively fine-tuned by the optimal contextualized observations sampling from its own outputs. Besides, to efficiently navigate in the exponentially large exploration space, NNetNav (Murty et al., 2024) utilizes the hierarchical structure of language instructions for fine-tuning, rendering the search process more manageable. It prunes interaction episodes automatically when an intermediate trajectory cannot be meaningfully annotated, optimizing exploration effectiveness. Besides, WebGUM (Furuta et al., 2023a) proposes a data-driven offline supervised fine-tuning process that adapts WebAgents for web navigation tasks based on instruction-following format problems and chain-of-thought examples across a variety of domains. The web navigation action output is in free-form text, enhancing the flexibility and adaptability to tackle real-world web navigation. In addition, LLMPA (Guan et al., 2023) implements an end-to-end finetuning framework to accomplish multi-step operations based on high-level user command by integrating various auxiliary modules, like instruction decomposition and action prediction. Notably, apart from these key supporting modules, LLMPA employs Controllable Calibration to meticulously evaluate predicted actions, effectively mitigating the hallucination issues prevalent in LLMs and enhancing overall execution reliability.
@@ 5 Trustworthy WebAgents
In recent years, the rapid evolution of LFM-empowered agents has ushered in a new era, profoundly reshaping various facets of web applications, including E-commerce, healthcare, and education. However, alongside this remarkable progress, a growing body of research and real-world incidents has highlighted the inherent risks and challenges associated with WebAgents (Yang et al., 2025; Schwartz et al., 2023; Deng et al., 2024). Recent studies have begun to reveal the potential threats posed by agents closely integrated with web systems. These threats range from unreliable and opaque decision-making in safety-critical contexts to the perpetuation of bias and unfairness towards marginalized communities (Deng et al., 2024; Li et al., 2024a). Moreover, concerns have been raised about the inadvertent exposure of user privacy and sensitive business information, further emphasizing the critical need to ensure the security of WebAgent technologies. Additionally, limitations in generalizability pose serious risks for WebAgents when confronted with diverse and unforeseen situations, potentially leading to critical failures in handling out-of-distribution data or operating across different domains. Consequently, developing advanced techniques for trustworthy WebAgents has become an area of increasing interest for researchers, primarily focusing on creating and deploying AI systems that prioritize Safety & Robustness, Privacy, and Generalizability.